OpenStudy (anonymous):

For the equation, y=-x+4 tell whether its graph passes through the first quadrant. Explain how you know.

OpenStudy (anonymous):

@zepdrix @blahhedup101 can you guys help me?

OpenStudy (anonymous):

Hey texas, thank you for helping me. I needed help on other questions and looked it up. I see that you also helped other students with a lot of questions :)

OpenStudy (texaschic101):

well....let me see y = -x + 4 In y = mx + b, the slope is m and the y intercept is b. so the slope is -1 and the y intercept is (0,4) the x intercept can be found by subbing in 0 for y 0 = -x + 4 x = 4....x intercept is (4,0) so start at y intercept (0,4) and come down 1 and to the right 1, then down 1, then to the right 1, and you should cross the x axis at (4,0) Once graphed, you will see that IT PASSES THROUGH the 1st quadrant

OpenStudy (anonymous):

Hmm, I am confused even if x is poistive it has a negative on the right

OpenStudy (texaschic101):

oh...you mean for the x intercept ? y = -x + 4 ....sub in 0 for y 0 = -x + 4 -- add x to both sides x = -x + x + 4 -- the x's on the right side cancel x = 4....the x intercept is (4,0) Is that what you are referring to ?

OpenStudy (texaschic101):

you can also check it by taking a point in quadrant 1 and checking....(2,2) is in quadrant 1. Lets sub y = -x + 4 2 = -2 + 4 2 = 2 (correct)

OpenStudy (anonymous):

Yes I understand. So the two answers it could be our Yes, for x>0 some y values are positive and yes for x<0 some y values are positive

OpenStudy (anonymous):

So it would be yes, for x>0 some y values are positive, right?

OpenStudy (texaschic101):

yes...x > 0 and some y values are positive

OpenStudy (anonymous):

ok thanks... I need help on another question... Do you think you can help me?

OpenStudy (texaschic101):

I can try

OpenStudy (anonymous):

Ok, thank you so much. \[km \div 3-2x=4\]

OpenStudy (anonymous):

Km/3-2x=4

OpenStudy (anonymous):

solve for x.

OpenStudy (texaschic101):

Km/3 - 2x = 4 -- start by multiplying the whole equation by 3 to get rid of the fraction giving us : Km - 6x = 12 -- subtract Km from both sides -6x = -Km + 12 -- now divide both sides by -6 x = 1/6Km - 2 sorry it took so long...got a little confused. But I finally got it

OpenStudy (anonymous):

Thank you so much! I was getting -6+km as the answer and it was wrong. I realised that I did not multiply the whole left side by 3 and only the denominator... Thank you so much!

OpenStudy (anonymous):

You are restoring faith in humanity by helping a 13 year old :)

OpenStudy (texaschic101):

sure thing :)

OpenStudy (texaschic101):

lol....glad to help